OpenStudy (anonymous):

Multiply (5√6-10√5)(4√6+8√5) I got so far 5(√6-2√5)4(√6+2√5) but am stuck here please help

OpenStudy (hoblos):

=5(√6-2√5)4(√6+2√5) =20(√6-2√5)(√6+2√5) difference of two squares =20[ (√6)^2 - (2√5)^2 ] =20[ 6 - 20] = 20(-14) =-280

OpenStudy (anonymous):

oh i see now. Ok thanks. have a few problems like this but i can handle them from here thanks for helping me understand

OpenStudy (hoblos):

my pleasure XD you are welcome at any time :)